OpenStudy (anonymous):

HELP PLEASE! I need to know how to do this...I will medal 4 questions on "Graphs to Solve Systems " can anyone help me please?

OpenStudy (anonymous):

What is the solution to the system of equations represented by the two equations? y=2/3x y=-2/3x+4 (blank, blank)